the time for a major exam to be completed is normally distributed with an average of 55 minutes and a standard deviation of 9 minutes. If 92% of the student completed the exam, when should the test be terminated?


1
Expert's answer
2021-02-28T15:49:21-0500

Given,

Average time "x=55 min"

Standard deviation "\sigma=9min"


Probability os student who completed the exam "=0.92"

As the exam time is normally distributed then The probabilty Who do not complete their exam

"= 1-P(x<0.92)=1-0.92=0.08"


The value of z at 0.08 from the standard table is -1.4


Using normal distribution formula,

"z=\dfrac{x-a}{\sigma}"


"\Rightarrow -1.4\times 9=55-a\\\Rightarrow a=55+12.6=67.6\text{ min}".


The test Should be terminated at 67.6 min.